OpenStudy (anonymous):

You invest an initial $1,000 in an account that has an annual interest rate of 2.5%, compounded annually. How much money will you have in the account after 6 years? Round your answer to the nearest whole number.

OpenStudy (xapproachesinfinity):

P=A(1+r)^t you need to use this A=1000 r=2.5/100 t=6